OpenStudy (anonymous):

Help, please. I need an explanation on how a problem like this would work... Joselyn is a manager at a sign-painting company. She has two painters, Allen and Brianne. Allen can complete a large project in 16 hours. Brianne can complete the project in 18 hours. Joselyn wants to know how long it will take them to complete the project together.

OpenStudy (anonymous):

\[\frac{ 1 }{ 16 }+\frac{ 1 }{ 18 }=\frac{ 1 }{ x }\] Is this the proper setup for a problem like this?